# Kac–Moody Symmetric Spaces: arbitary symmetrizable complex or almost split real type

### v3

@@ 3.6 The Iwasawa Decomposition
where $C\in\mathcal{C}_{\pm}$ denotes a chamber in $\Delta_{\pm}$, $w\in\mathcal{W}$ an element of the Weyl group and $\delta_{\pm}$ the Weyl distance in relation to the two halves of the twin building. By the previous discussion $K_{l}^{\pm}$ acts chamber-transitive on
$\bigcup_{w\in\mathcal{W}\,,\;l(w)\leq l}E_{w}(C).$
Moreover, this is a finite union of compact sets, see [HKM13, Corollary 3.10].
Using again the chamber-transitivity it follows that we can rewrite the union as the quotient
$K_{l}^{\pm}/(K_{l}^{\pm}\cap B_{\pm}).$
Recall that $M$ is the compact subgroup of $T$, in particular $M$ is a Lie group, and that $K_{l}^{\pm}\cap B_{\pm}\subseteq M$. Thus, by the discussion around [Pal61, Corollary 4.1], it follows that
@@
are fibers. $K_{l}^{\pm}$ are closed sets so it follows that
$\pi\colon K_{l}^{\pm}\to K_{l}^{\pm}/(K_{l}^{\pm}\cap B_{\pm})$
are also fibers and the corresponding fiber bundles are compact. Thus $K_{l}^{\pm}$ is a compact set for every $l\in\mathbb{N}$, therefore, the pair $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$ pair. Together with the fact that
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is non-spherical. Then each of the multiplication maps
@@
which are by construction inverse to $m_{p}m$ and they are continuous, since $G$ is a topological group and $\iota_{\pm}$ is a homeomorphisms. It is left to check, that the multiplication maps $A\times U_{\pm}\to AU_{\pm}$ are open, but this follows from [HKM13, Proposition 7.29]. ∎
A useful application of the Iwasawa decomposition is the following lemma, which can be found for centered Kac–Moody groups over $\mathbb{R}$ in [FHHK20, Lemma 3.27].
@@
The following applies: $\tau(t)=A$ and $T\subseteq N_{G}(T)$, i.e. $A\subseteq N_{G}(T)\cap\tau(G)$. For the converse implication let $g\in N_{G}(T)\cap\tau(G)$. By 3.15 there exists elements $u\in U_{+}$, $t\in T$ and $k\in K$, such that
@@
Since $g\in N_{G}(T)$ it follows that $g=\tau(t)$ and hence $g\in A$. ∎
@@ 4.1 A criterion for symmetric spaces
In the following section we prove a statement in the context of buildings that allows us to show that $K\cap\tau(G)=\{e\}$. This result will be a very powerful statement to show that a Kac–Moody group can be transformed into a symmetric space in the sense of Loos. The following lemma is very technical and uses building theory arguments, for a brief introduction see Section 3.3. Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\mathbb{A}$ be a symmetrizable, indecomposable generalized Cartan matrix.
In order to state the lemma, call an element $g\in G$ of a Kac–Moody group $G$ with the property
$\Theta(g)=g^{-1}$
symmetric.
@@
Let $G$ be a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$ and denote with $\Delta_{\pm}$ the associated twin building. For a symmetric element $g\in G$ the statements below are equivalent:
@@
    The element $g$ fixes a $\Theta$-stable twin apartment chamberwise. $\Theta$-stable means that a twin apartment of $\Delta_{\pm}$ is invariant as a set under the action of $\Theta$.
@@
    Hence one obtains by a straight forward calculation,
    $\Theta(D)=\Theta(g.D)=g^{-1}.\Theta(D),$
    that $g$ stabilizes the $\Theta$-stable twin apartment $\Sigma(D,\Theta(D))$, since $\Theta(D)$ is opposite of $D$.
@@
    Let $D^{\prime}\in\Delta_{+}$ be a chamber in the positive half of the twin building which is stabilized by $g$, also let $D\in\Delta_{+}$ be an arbitrary chamber and $n\in\mathbb{Z}$. Then we know that
@@
    Now let $R$ be a spherical residue stabilized by $g$ and by the computation
@@
    and the previous comment, follows that $\Theta(R)$ is also stabilized by $g$. Hence the symmetric element $g$ stabilizes the spherical residue in each half of $\Delta_{\pm}$.
@@
    By [AB08, Proposition 67.27], one knows that the stabilizer of a spherical residue of type $J$ equals a spherical parabolic subgroup $P_{J}$, compare this with the discussion before [Mar18, Proposition 7.75]. Since $g$ stabilizes $R$ and an opposite residue $\Theta(R)$, it is an element of
@@
    Using the fact, that a parabolic subgroup, in the context of a Kac–Moody group, has a Levi decomposition (see [Rém02, 6.2.2 (Théorème)] or [Mar18, Proposition 7.75]) one can analyze this intersection. Therefore, denote by
@@
    the Levi decomposition of $P_{J}$ and note that $\Theta(P_{J})$ is obtained by applying the Chevalley involution.
@@
    Since $g$ is an element of the intersection, it follows that it can be written in two ways:
@@
    since $y^{-1}$ is an element of the stabilizer. It yields that $u_{-}\in P_{J}\cap U_{J}^{-}=\{e\}$, and analogously one can show with $\Theta(P_{J})$ that $u_{+}$ must also be trivial.
@@
    The Chevalley involution translates for these matrices to transpose-inverse, composed with complex conjugation if $\mathbb{K}=\mathbb{C}$. Taking a closer look on the property that $g$ is symmetric turns $M_{g}$ it into a Hermitian matrix, in particular
@@
    Thus, $M_{g}$ is diagonalizable and hence it is conjugated to an diagonal matrix $D_{g}$. Translated back into the language of Kac–Moody groups, this leads to the fact that $g$ is conjugate to an element of the standard torus $T$ of the Kac–Moody group $G$. Since the standard torus fixes the pair of fundamental chambers $(C_{+},C_{-})$ in $\Delta_{\pm}$, a conjugated element fixes a translated pair of the fundamental chambers. In detail if $T$ fixes $C_{\pm}$ and $aga^{-1}\in T$ denotes a conjugated element in $T$, one calculates
@@
    By the argument from above, $g$ fixes then the twin apartment $\Sigma(a^{-1}.C_{\pm},a^{-1}.\Theta(C_{\pm}))$ chamberwise. Note that all of these arguments apply to both halves of the twin building for reasons of symmetry.
The following lemma is a consequence of the latter statement and allows us to prove that a special quotient of a Kac–Moody group is a symmetric space. Thus, the lemma is one of the central parts of the theory we develop in this article.
The version for non-affine centered Kac–Moody groups over $\mathbb{R}$ can be found at [FHHK20, Proposition 3.38].
@@
Let $G$ be a Kac–Moody group and denote by $\Delta_{\pm}$ the associated twin building. Further, let $g$ be an element of $\tau(G)$, i.e. $g=h\Theta(h)^{-1}$, for an element $h\in G$. Then one calculates
@@
Thus, if $g\in K\cap\tau(G)$, one realizes that $g=\Theta(g)=g^{-1}$, from which it follows that $g$ is symmetric and has order at most $2$. This leads to the fact that the orbits of $g$ are bounded as in point 4 of 4.1. Hence by the proof of 1 from 4.1 one knows, that $g$ fixes the $\Theta$-stable twin apartment $\Sigma(C,\Theta(C))$ chamberwise. Thus $g$ is contained in a corresponding torus $T^{\prime}$ of $G$. By 3.15 it follows, that $K$ acts transitively on both halves of the associated twin building $\Delta_{\pm}$.
Hence there is a $k\in K$ such that $kT^{\prime}k^{-1}=T$. It follows that the element
@@ 4.2 Kac–Moody Symmetric Space
Note that this definition of the reflection map $\mu$ also holds for any topological group with an involution. Moreover, the first three properties in 2.1 can be proved without any assumption about the group, see the calculation in the proof of [FHHK20, Proposition 4.2]. Thus any topological group with an involution that satisfies the last property gives rise to a symmetric space.
@@
    To check if $\tau(G)\cap K$ is trivial, we use that $\tau(A)$ can be diagonalized: let $S$ be a orthogonal real matrix such that
@@
    and it follows that the entries of $D$ are ±$\pm 1$. But by the positive semi-definite property only $+1$ is possible.
@@ 4.3 The Group Model
If $g\in G$ is an element of the kernel, then $ghK=hK$ for all $h\in G$, in particular for $h\in K$ follows $ghK=K$ which implies that $g\in K$. Now assume that $g\in K$ acts trivially on $hK$ for an element $h\in G$. Then we can compute $ghK=hK\Leftrightarrow h^{-1}gh\in K$ and by using 3.14 (A1) it follows that $\tau(h^{-1}gh)=e$. By a direct calculation this leads to $g\tau(h)g^{-1}=\tau(h)$, and thus the kernel of the action is given by $C_{K}(\tau(G))$.
@@
This leads to the observation, that the map
$\zeta\colon\tau(G)\to T(G/K,eK)$
is surjective. Furthermore, if an element is contained in the kernel, it must be contained in $K$ by the previous calculation. Since $\tau(G)\cap K=\{e\}$, see 4.3, it follows that the kernel of $\zeta$ is trivial. By transport of structure, the map establishes a symmetric space structure on $\tau(G)$, where the reflection map is given by $\mu_{\tau}(g,h)=gh^{-1}g$.
@@
On the other hand, one knows that $G^{\dagger}=\langle\tau(G^{\dagger})\rangle$ (cp. [FHHK20, Proposition 3.39]) as well as $\tau(T)=A$. Note that [FHHK20, Proposition 3.39] is formulated for centered Kac–Moody groups with an RGD system, and $G^{\dagger}$ is a centered Kac–Moody group with an RGD system, see the discussion after 3.13. Thus
@@
Further, $\tau(h)^{-1}=\tau(\Theta(h))$, hence the transvection group is generated by elements of the form

### v4

@@ 3.6 The Iwasawa Decomposition
where $C\in\mathcal{C}_{\pm}$ denotes a chamber in $\Delta_{\pm}$, $w\in\mathcal{W}$ an element of the Weyl group and $\delta_{\pm}$ the Weyl distance in relation to the two halves of the twin building. The above isomorphism of $K$-sets between $\Delta_{\pm}$ and $K/(K\cap B_{\pm})=K/(K\cap T)$ induces a bijection between
$\bigcup_{w\in\mathcal{W}\,,\;l(w)\leq l}E_{w}(C)\quad\quad\text{and}\quad\quad K_{l}^{\pm}/(K_{l}^{\pm}\cap B_{\pm}).$
This is a finite union of compact sets, see [HKM13, Corollary 3.10]. Recall also that the finite group $M$ is the maximal compact subgroup of $T$ and that $K_{l}^{\pm}\cap B_{\pm}\subseteq M$. Thus
@@
is a finite covering, and so are
$\pi\colon K_{l}^{\pm}\to K_{l}^{\pm}/(K_{l}^{\pm}\cap B_{\pm}).$
Thus $K_{l}^{\pm}$ is a compact set for every $l\in\mathbb{N}$, therefore, the pair $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$-pair. Together with the fact that
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is a symmetrizable, irreducible, non-spherical Cartan matrix. Then each of the multiplication maps
@@
which are by construction inverse to $m_{\pm}$ and they are continuous, since $G$ is a topological group and $\iota_{\pm}$ is a homeomorphisms. It remains to check, that the multiplication maps $A\times U_{\pm}\to AU_{\pm}$ are open, but this follows from [HKM13, Corollary 7.30], as the Levi decomposition is a semi-direct product of closed subgroups. ∎
A useful application of the algebraic Iwasawa decomposition is the following lemma, which can be found for centered Kac–Moody groups over $\mathbb{R}$ in [FHHK20, Lemma 3.27].
@@
The following applies: $\tau(T)=A$ and $T\subseteq N_{G}(T)$, i.e. $A\subseteq N_{G}(T)\cap\tau(G)$. For the converse implication let $g\in N_{G}(T)\cap\tau(G)$. By 3.15 there exist elements $u\in U_{+}$, $t\in T$ and $k\in K$, such that
@@
Since $g\in N_{G}(T)$ we conclude that $g=\tau(t)$ and hence $g\in A$. ∎
@@ 4.1 A criterion for symmetric spaces
We start with a technical building-theoretic lemma; for a brief introduction to buildings see Section 3.3. Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\mathbb{A}$ be a symmetrizable, indecomposable generalized Cartan matrix.
An element $g\in G$ of a Kac–Moody group $G$ is called symmetric, if
$\Theta(g)=g^{-1}.$
Further we call a twin apartment $\Theta$-stable, if the to $G$ associated twin building is invariant as a set under the action of $\Theta$. Recall that $\Delta_{\pm}=G/B_{\pm}$, then we consider the action of $\Theta$ on the building as follows
$\Theta(g.B_{\pm})=\Theta(g).B_{\mp}.$
@@
Let $G$ be a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$ and denote by $\Delta_{\pm}$ the associated twin building. For a symmetric element $g\in G$ the statements below are equivalent:
@@
    The element $g$ fixes a $\Theta$-stable twin apartment chamberwise.
@@
In the following we use the notation of a twin apartment given in [AB08, Definition 5.176]: Let $C$ and $C^{\prime}$ be opposite chambers, i.e. $C\in\mathcal{C}_{\varepsilon}$, where $\varepsilon\in\{+,-\}$ and $C^{\prime}\in\mathcal{C}_{-\varepsilon}$. Then $\Sigma(C,C^{\prime})$ is defined as the set
$\Sigma(C,C^{\prime})\coloneqq\{D\in\mathcal{C}_{\varepsilon}\mid\delta_{\varepsilon}(C,D)=\delta^{\star}(C^{\prime},D)\}.$
For details on the notation, see Section 3.3 or [AB08, Chapter 5.8].
Furthermore, a $J$-residue is a certain equivalence class of chambers: two chambers $C,D$ in $\Delta_{\varepsilon}$ are called $J$-equivalent if $\delta_{\varepsilon}(C,D)\in W_{J}$, where $\mathcal{W}_{J}\coloneqq\langle J\rangle$ for a set $J\subseteq S$ w.r.t. Coxeter system $(\mathcal{W},S)$ of $\Delta_{\pm}$. For details see [AB08, Chapter 5.3].
@@
    Hence
    $\Theta(D)=\Theta(g.D)=g^{-1}.\Theta(D).$
    In particular, one concludes that $D=g.D$ implies $\Theta(D)=g.\Theta(D)$. Moreover, $g$ stabilizes the $\Theta$-stable twin apartment $\Sigma(D,\Theta(D))$, since $\Theta(D)$ is opposite of $D$.
@@
    Let $D^{\prime}\in\Delta_{+}$ be a chamber in the positive half of the twin building which is stabilized by $g$. This implies that $\Theta(D^{\prime})\in\Delta_{-}$ is stabilized by $g$ as well. Also let $D\in\Delta_{+}$ be an arbitrary chamber and $n\in\mathbb{Z}$. Then we know that
@@
    Now let $R$ be a spherical residue, i.e. a residue where $\mathcal{W}_{J}$ is finite, stabilized by $g$ and by the computation
@@
    and the previous comment, we have $\Theta(R)$ is also stabilized by $g$. Hence the symmetric element $g$ stabilizes the spherical residue in each half of $\Delta_{\pm}$.
@@
    By [AB08, Proposition 67.27], one knows that the stabilizer of a spherical residue of type $J$ equals a spherical parabolic subgroup $P_{J}$, compare this with the discussion before [Mar18, Proposition 7.75]. Since $g$ stabilizes $R$ and the opposite residue $\Theta(R)$, it is an element of
@@
    Using the fact, that a parabolic subgroup, in the context of a Kac–Moody group, has a Levi decomposition (see [Rém02, Theorem 6.2.2] or [Mar18, Proposition 7.75]) one can analyze this intersection. Therefore, denote by
@@
    the Levi decomposition of $P_{J}$ and note that $\Theta(P_{J})$ is obtained by applying the Chevalley involution. Note that in the convention of [Rém02, Chapter 6.2.2] $G_{J}$ is denoted by $M(F)$.
@@
    Since $g$ is an element of the intersection, we conclude that it can be written in two ways:
@@
    since $y^{-1}$ is an element of the stabilizer (see the second statement in the theorem of [Rém02, Chapter 6.2.2]). From the proof of [Rém02, Lemma 6.2.1] (and for the sake of understanding with the notation from [Rém02, Theorem 6.2.2]) we obtain that $u_{-}\in P_{J}\cap U_{J}^{-}=\{e\}$, and analogously one can show with $\Theta(P_{J})$ that $u_{+}$ must also be trivial.
@@
    For a more detailed discussion of this embedding we refer to [PR94, Chapter 3.2] and in particular to [PR94, Proposition 3.13]. There it is explained that a spherical Kac–Moody group (i.e. in this concrete case $G_{J}$) can be embedded in $\operatorname{GL}_{n}(\mathbb{K})$. This is done by using an appropriate conjugation so that the Chevalley involution is translated into the transpose inverse composed with complex conjugation when $\mathbb{K}=\mathbb{C}$. Note that this embedding, where the Chevalley involution is translated in a pleasant way, must be chosen and is not given or natural.
    Taking a closer look on the property that $g$ is symmetric turns $M_{g}$ it into a Hermitian matrix, in particular
@@
    Thus, $M_{g}$ is diagonalizable and hence it is conjugated to a diagonal matrix $D_{g}$. In particular, $g$ is conjugate to an element of a maximal torus $T_{J}$ of $G_{J}$. The torus $T_{J}$ embeds into a maximal torus of $G$, and so $g$ is conjugate to an element of the standard torus $T$ of the Kac–Moody group $G$. Since the standard torus fixes the pair of fundamental chambers $(C_{+},C_{-})$ in $\Delta_{\pm}$, a conjugated element fixes a translated pair of the fundamental chambers. In detail if $T$ fixes $C_{\pm}$ and $aga^{-1}\in T$ denotes a conjugated element in $T$, one calculates
@@
    This completes the proof of this implication.
Next we use the preceding lemma to conclude that the homogenenous space $G/K$ for a Kac–Moody group $G$ is a symmetric space. This lemma is one of the key ingredients of the theory we develop in this article.
The version for non-affine centered Kac–Moody groups over $\mathbb{R}$ is [FHHK20, Proposition 3.38].
@@
Denote by $\Delta_{\pm}$ the associated twin building to the Kac–Moody group $G$. Further, let $g$ be an element of $\tau(G)$, i.e. $g=h\Theta(h)^{-1}$, for an element $h\in G$. Then one calculates
@@
Thus, if $g\in K\cap\tau(G)$, one realizes that $g=\Theta(g)=g^{-1}$, from which we conclude that $g$ is symmetric and has order at most $2$. This leads to the fact that the orbits of $g$ are bounded as in point 4 of 4.1. Hence by the proof of 3 to 1 from 4.1 one knows, that $g$ fixes the $\Theta$-stable twin apartment $\Sigma(C,\Theta(C))$ chamberwise. Thus $g$ is contained in a corresponding torus $T^{\prime}$ of $G$. From 3.15 we know that the group $G$ can be decomposed as $G=KAU_{\pm}$. This can be turned into the decomposition
$G=KB_{\pm}$
by using $B_{\pm}=T\times U_{\pm}$ and the properties described in 3.14. This decomposition yields that $K$ acts transitively on both halves of the associated twin building $\Delta_{\pm}$.
Hence there is a $k\in K$ such that $kT^{\prime}k^{-1}=T$. Thus the element
@@ 4.2 Kac–Moody Symmetric Space
Note that this definition of the reflection map $\mu$ also holds for any topological group with an involution. Moreover, the first three properties in 2.1 can be proved without any assumption about the group, see the calculation in the proof of [FHHK20, Proposition 4.2]. Thus any topological group with an involution that satisfies the last property gives rise to a symmetric space. Here are two examples of such symmetric spaces.
@@
    To check if $\tau(G)\cap K$ is trivial, we use that $\tau(A)$ can be diagonalized: let $S$ be an orthogonal real matrix such that
@@
    and it can be deduced that the entries of $D$ are ±$\pm 1$. But by the positive semi-definite property only $+1$ is possible.
@@ 4.3 The Group Model
If $g\in G$ is an element of the kernel, then $ghK=hK$ for all $h\in G$. In particular, this assertion also applies if $h\in K$ and in this case $ghK=K$, which implies $g\in K$.
Let $g$ now be an element of $K$, which acts trivially on $hK$ for an element $h\in G$. Then we can compute $ghK=hK\Leftrightarrow h^{-1}gh\in K$ and by using 3.14 (A1) we have $\tau(h^{-1}gh)=e$. By a direct calculation this leads to $g\tau(h)g^{-1}=\tau(h)$, and thus the kernel of the action is contained in $C_{K}(\tau(G))\subseteq K$.
For the converse consider an element $c\in C_{K}(\tau(G))$ and recall the following properties: The group $G$ can be decomposed into $G^{\dagger}$ and $T$, where $G^{\dagger}$ is generated by the root subgroups of $G$, see 3.11. This allows us to apply [FHHK20, Proposition 3.39], which gives us the following
$\left\langle\tau\left(G^{\dagger}\right)\right\rangle=G^{\dagger}.$
Now let $c\in C_{K}(\tau(G))$ and $h=tg\in G=TG^{\dagger}$ the we can calculate
$chK=ctgK=tcgK=tgcK=hcK=hK.$
We have therefore shown the reverse inclusion that $C_{K}(\tau(G))$ is contained in the kernel.
@@
From this, one can define a map $\zeta\colon\tau(G)\to T(G/K,eK)$ by sending an element $\tau(g)$ of $\tau(G)$ to the coset $\tau(g)C_{K}(\tau(G))$. Then the previous observation shows that the map $\zeta$ is surjective. Furthermore, if an element is contained in the kernel of $\zeta$, it must be contained in $K$ by the previous calculation. Since $\tau(G)\cap K=\{e\}$, see 4.3, we conclude that the kernel of $\zeta$ is trivial. By transport of structure, the map establishes a symmetric space structure on $\tau(G)$, where the reflection map is given by $\mu_{\tau}(g,h)=gh^{-1}g$.
@@
Recall that $G^{\dagger}$ is the derived subgroup of $G$, see the proof of 3.12. Note that eff in the index stands for effectively, which means that the group acts faithfully on $G/K$.
@@
On the other hand, one knows that $G^{\dagger}=\langle\tau(G^{\dagger})\rangle$ (cf. [FHHK20, Proposition 3.39]) as well as $\tau(T)=A$. Note that [FHHK20, Proposition 3.39] is formulated for centered Kac–Moody groups with an RGD system, and $G^{\dagger}$ is a centered Kac–Moody group with an RGD system, see the discussion after 3.13. Thus
@@
Further, $\tau(h)^{-1}=\tau(\Theta(h))$, hence the transvection group is generated by elements contained in